[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jun 28, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jun 28 09:26:11 UTC 2017


HSU Shift Log for Wed Jun 28 2017 05:26:05 GMT-0400 (EDT)

Shift: A

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: GOES-16 data stopped updating in NAWIPS and AWIPS  
between 01Z and 02Z

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170628/919d5f53/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list